The measures of the angles of a triangle in degrees can be expressed by the ratio 5:6:7. What is the sum of the measures of the two larger angles?
Answer by stanbon(75887)

The measures of the angles of a triangle in degrees can be expressed by the ratio 5:6:7. What is the sum of the measures of the two larger angles?
-----
Equation:
5x + 6x + 7x = 180
18x = 180
x = 10
-----
6x + 7x = 13x = 13*10 = 130 degrees
==================
Cheers,
Stan H.
==================